OpenStudy (anonymous):

please solve this Equation. (x-4)/(x+2)=5

OpenStudy (rina.r):

x2 - 2x/-4x-8=5 -2x / 32 =5 -16x =5 x=-16/5 =3.2 i think..

OpenStudy (anonymous):

Multiply both sides by x+2: x-4=5(x+2) Distribute the 5: x-4=5x+10 Subtract 5x from both sides. Add 4 to both sides: -4x=14 Divide both sides by -4: x=-7/2 Hope that helps

OpenStudy (anonymous):

Thanks a lot.